OpenStudy (anonymous):

A 5.0 m portion of wire carries a current of 4.0 A from east to west. It experiences a magnetic field of 6.0 × 10^-4 T running from south to north. What is the magnitude and direction of the magnetic force on the wire? 1.2 × 10^-2 N downward 1.2 × 10^-2 N upward 2.4 × 10^-2 N downward 2.4 × 10^-2 N upward

OpenStudy (irishboy123):

you should know all of this this http://hyperphysics.phy-astr.gsu.edu/hbase/magnetic/forwir2.html in terms of magnitude, \(F = B\times I \times L\), so just plug in the numbers in terms of direction, you have \( \hat{ I}_{E2W} \times \hat{B}_{S2N} = \hat{F}_{out \ of \ page}\) for that you need the right hand rule, see attached.